The left and right page numbers of an open book are two consecutive integers whose sum is 359. Find these page numbers
Answer by TimothyLamb(4379)

y = x + 1
x + y = 359
---
put the system of linear equations into standard form
---
x - y = -1
x + y = 359
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 179
y = 180
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php